OpenStudy (anonymous):

Mr.gomez paid a total of $267 for the repairs on his car. the cost of the labor was two thirds of the total charge. Find the charge for the labor that was done.

OpenStudy (lilg132):

178 267 / 3 = 89 89 * 2 = $178 remember to click good answer

OpenStudy (anonymous):

$267 * 2/3 =cost of labor =$178